OpenStudy (anonymous):

If the Discriminant of an equation is negative, which of the following is true of the equation

OpenStudy (anonymous):

if this is quadratic then discriminant >0 then two real roots discriminnt = 0 1 real root discriminant < 0 complex roots

OpenStudy (anonymous):

which two values of x are roots of the polynomial below x^2 + 5x + 9

OpenStudy (anonymous):

you can do this in two ways 1) Sick it into the quadratic formula.\[x= (-b \pm \sqrt{b ^{2} - 4a2}) \div 2a\] 2) try to factorise where you wor out all the factors of nine and the two that add to +5 is the anwer

OpenStudy (anonymous):

Ill show you

OpenStudy (anonymous):

\[(-5+\sqrt{-11})\div2 \] \[(-5-\sqrt{-11})\div2 \] because 5^2 - 4x1x9 is negative so it has no real roots

OpenStudy (anonymous):

hmm the quadratic eqn is wrong should be :\[(-b \pm \sqrt{b^2-4 \times a \times c})\div 2 \times a\]